# DMUS-MP-25/04

### v1

@@ 3 Relaxing the special coupling
The basic idea is now to let $q$ be generic in this reformulation of the problem, and see where this leads. The hope is that this introduces the mixed flux in the $AdS_{3}$ case - we mean this purely in terms of functional form of the $S$-matrix, since it is clear that the spectra are different [23] and it is not immediate to us to imagine how they might smoothly deform one into another. So, we could simply recalculate the $S$-matrix with generic $q$, since we can just “forget” that $q$ was special in the above formulas. Applying this to Fendley-Intriligator of course returns back the ordinary Sine-Gordon. For $AdS_{3}$ the result, although potentially interesting per se, is not immediately connected to mixed-flux scattering: this is due to the fact that the mixed-flux relativistic coproduct is twisted in a very particular way [35, 23], so we have to work a bit harder.
There is in fact a way to obtain exactly the mixed flux scattering matrix using this philosophy, by adjusting things in such a way that the deformation of the coproduct is precisely the one observed in the mixed flux case. It works like this. We could have taken the following alternative route to the pure RR massless $AdS_{3}$ $S$-matrix:
$S_{k}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&e^{\frac{\theta}{2}}\big{(}1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big{)}&\epsilon(\theta)&0\\ 0&e^{i\frac{\pi}{k}}\big{(}1-e^{\theta}+e^{i\frac{\pi}{k}+\theta}\epsilon(\theta)\big{)}&e^{\frac{\theta}{2}}\big{(}1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big{)}&0\\ 0&0&0&e^{\theta}-e^{i\frac{\pi}{k}}\big{(}1+e^{\theta}\big{)}\epsilon(\theta))\end{pmatrix}$
$\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}e^{i\pi(q+i)}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}e^{-i\pi(q+i)}\otimes\hat{F}$,
$\qquad|\phi|=|\psi|=0,\quad|\hat{E}|=|\hat{F}|=0,\quad q_{0}=-i,\qquad q=\frac{1}{k}-i,\quad\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi},\quad\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi},$
@@
$\epsilon(\theta)=\frac{1}{e^{\frac{i\pi}{k}}+\frac{1}{e^{g}-e^{g+\theta}}},\qquad g\,\,\mbox{constant}.$ (3.1)
We have solved YBE by noticing that it reduces for a single condition for the function $\epsilon(\theta)$ with various arguments. Taking repeated derivatives w.r.t. $\theta_{i}$, $i=1,2,3$, and setting to zero these derivatives, eventually delivers the unique solution (3.1) with a constant of integration remaining arbitrary.
@@
$S_{k}\to\lim\limits_{k\to\infty,\epsilon(\theta)\to\tanh\frac{\theta}{2}}\to\,\,\,\,\qquad\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&\tanh\frac{\theta}{2}&0\\ 0&-\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (3.2)
$\Delta{\hat{E}}\to\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}\otimes\hat{E},\qquad\Delta{\hat{F}}\to\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}\otimes\hat{F}$.
The choice displayed in the limit (3.2) consistently corresponds to setting $k\to\infty$, and then setting $e^{g}=-\frac{1}{2}$, in (3.1).
@@
$S_{k}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&e^{\frac{\theta}{2}}\big{(}1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big{)}&\epsilon(\theta)&0\\ 0&e^{i\frac{\pi}{k}}\big{(}1-e^{\theta}+e^{i\frac{\pi}{k}+\theta}\epsilon(\theta)\big{)}&e^{\frac{\theta}{2}}\big{(}1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big{)}&0\\ 0&0&0&e^{\theta}-e^{i\frac{\pi}{k}}\big{(}1+e^{\theta}\big{)}\epsilon(\theta))\end{pmatrix}$
$\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}e^{i\pi(q+i)}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}e^{-i\pi(q+i)}\otimes\hat{F}$,
$\qquad|\phi|=|\psi|=0,\quad|\hat{E}|=|\hat{F}|=0,\quad q_{0}=-i,\qquad q=\frac{1}{k}-i,\quad\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi},\quad\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi},$
$\qquad\qquad\to\lim\limits_{\epsilon(\theta)\to-i\frac{\sinh\frac{\theta}{2}}{\sin\big{(}\frac{\pi}{k}-\frac{i\theta}{2}\big{)}}}\to\,\,\,\,\qquad\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{i\sin\frac{\pi}{k}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&\frac{\sinh\frac{\theta}{2}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&0\\ 0&\frac{\sinh\frac{\theta}{2}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&\frac{i\sin\frac{\pi}{k}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&0\\ 0&0&0&\frac{\sinh\big{(}\frac{i\pi}{k}-\frac{\theta}{2}\big{)}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}.\end{pmatrix}$ (3.3)
The limit coincides with the left-left relativistic mixed-flux scattering - see for instance [23]. The choice displayed in the limit (3.3) consistently corresponds to setting setting $e^{g}=\frac{i}{2}\mbox{csc}\frac{\pi}{k}$, in (3.1).
@@
$S_{k}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{\theta}{2}}}{1-e^{i\frac{\pi}{k}+g}\big{(}e^{\theta}-1\big{)}}&\frac{1}{e^{\frac{i\pi}{k}}+\frac{1}{e^{g}-e^{g+\theta}}}&0\\ 0&\frac{e^{i\frac{\pi}{k}}\big{(}1-e^{\theta}\big{)}\big{(}1+e^{i\frac{\pi}{k}+g}\big{)}}{1-e^{i\frac{\pi}{k}+g}\big{(}e^{\theta}-1\big{)}}&\frac{e^{\frac{\theta}{2}}}{1-e^{i\frac{\pi}{k}+g}\big{(}e^{\theta}-1\big{)}}&0\\ 0&0&0&\frac{1+e^{\theta}}{1-e^{i\frac{\pi}{k}+g}\big{(}e^{\theta}-1\big{)}}-1\end{pmatrix}$ (3.4)
pure RRinterpolationmixed flux
@@
$k\to\infty,\,\,e^{g}=-\frac{1}{2},\qquad\qquad\qquad\qquad\qquad k\,\,\mbox{generic},\,\,\,e^{g}=\frac{i}{2}\mbox{csc}\frac{\pi}{k}.$ (3.5)
We have summarised the situation in fig. 3. Notice also that $S_{k}$ satisfies braiding unitarity provided $\Phi(\theta)\Phi(-\theta)=1$. It is also physical-unitary if $\Phi(\theta)$ is a pure phase for real $\theta$ and real $k$, and if a certain condition holds on $g$, specifically
$\mbox{Im}(g)=-i\log\Big{[}-\frac{1}{2}e^{-\mbox{Re}(g)-\frac{i\pi}{k}}\Big{(}1+\sqrt{1-4e^{2\mbox{Re}(g)}}\Big{)}\Big{]},$ (3.6)
Figure 3: $S_{k}$ interpolates between pure RR massless $AdS_{3}$ and relativistic mixed flux.
In mathematical terms, we have basically introduced a Drinfeld twist, which can be formalised for example as:
$R_{k}\to\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}e^{i\pi(q+i)}\otimes\hat{E}=T(\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}\otimes\hat{E})T^{-1}\leftarrow T\bar{R}T^{-1},$
$\qquad\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}e^{-i\pi(q+i)}\otimes\hat{F}=T(\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}\otimes\hat{F})T^{-1}$
$\qquad T=e^{i\frac{\pi(q+i)}{2}\mathbbmss{1}\otimes\sigma_{3}},\qquad q=\frac{1}{k}-i,$ (3.7)
where $\sigma_{3}=E_{11}-E_{22}$. From the general theory of Drinfeld twists one obtains the relation
$R_{k}=PS_{k},\qquad P=E_{ij}\otimes E_{ji},\qquad T_{21}^{-1}R_{k}T=\bar{R}.$ (3.8)
Similar considerations are at the basis of [54], see also [4] and section 4.3 of [32] for example. Notice that one can use either the $S$-matrix or the associated $R$-matrix - we actually use the appropriate $R$-matrices throughout the paper to prove the Yang-Baxter equation.
3.1 Fractional route
It is again interest to obtain an $S$-matrix which depends on a (more general) unspecified fractional fermionic in such a way to interpolate between Sine-Gordon and $AdS_{3}$ when q=iq=-i, but then to obtain the mixed flux $S$-matrix in another regime.
More general fractional interpolation:
$S_{a+}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{i\pi(\alpha+\beta)}{2}}\big{(}-1+e^{\frac{\theta}{2}}\kappa(\theta)\big{)}}{\sigma}&\kappa(\theta)&0\\ 0&\kappa(\theta)&\sigma e^{\frac{-i\pi(\alpha+\beta)}{2}}\big{(}-1+e^{-\frac{\theta}{2}}\kappa[\theta]\big{)}&0\\ 0&0&0&-1+2\kappa(\theta)\cosh\frac{\theta}{2}\end{pmatrix}$
subject to
$\sigma^{2}=1,\qquad$ (3.9)
$\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+e^{-2H(\log[q]-\log[-i])+\mbox{Fracto}}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+e^{2H(\log[q]-\log[-i])-\mbox{Fracto}}\otimes\hat{F}$,
$|\phi|=0,\qquad|\psi|=0,\quad|\hat{E}|=0,\quad|\hat{F}|=0,\qquad\mbox{Fracto}=\begin{pmatrix}\alpha&0\\ 0&\beta-1\end{pmatrix},\quad,\quad\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi}\quad\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi},$
@@ 3.1 Fractional route
$e^{i\pi\alpha}=-q^{2}\,e^{i\pi\beta},\qquad q=i\sigma e^{\frac{i\pi}{2}(\alpha-\beta)}.$ (3.10)
@@
[…]
@@
$S_{a+}\to\mbox{massless $AdS_{3}$ \ \ \ \ \ \ \ if}\qquad q\to-i,\,\,\alpha,\beta\to 0,\qquad\kappa(\theta)=\mbox{sech}\frac{\theta}{2},\qquad\sigma=-1.$ (3.12)
@@
$\sigma=-e^{\frac{i\pi(\alpha+\beta)}{2}-\frac{i\pi}{k}},\qquad\kappa(\theta)=\frac{e^{\frac{\theta}{2}}\big{(}e^{\frac{2i\pi}{k}}\ -1\big{)}}{e^{\frac{2i\pi}{k}+\theta}-1},$ (3.13)
@@
$S_{a+}\to\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{i\sin\frac{\pi}{k}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&\frac{\sinh\frac{\theta}{2}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&0\\ 0&\frac{\sinh\frac{\theta}{2}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&\frac{i\sin\frac{\pi}{k}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&0\\ 0&0&0&\frac{\sinh\big{(}\frac{i\pi}{k}-\frac{\theta}{2}\big{)}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}\end{pmatrix}$, (3.14)
exactly the mixed flux result. Given that $\sigma$ is a sign, we must have
@@
$\alpha=\beta=(\frac{1}{k}+2n+1,\frac{1}{k}+2n),\qquad\qquad n\in\mathbbmss${Z}, (3.15)
@@
$|\phi|=\frac{1}{k}+1,\qquad|\psi|=\frac{1}{k},$ (3.16)
@@
$|\phi|=\frac{1}{k},\qquad|\psi|=\frac{1}{k}-1.$ (3.17)
@@
$\kappa(\theta)=\frac{e^{\frac{\theta}{2}}}{1+2\,d\,e^{\frac{\theta}{2}}\sinh\frac{\theta}{2}},\qquad d\,\ \mbox{constant}.$ (3.18)
@@
$S_{a+}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{\theta}{2}}}{1+d(e^{\theta}-1)}&\frac{(1-d)e^{\frac{i\pi(\alpha+\beta)}{2}}\big{(}e^{\theta}-1\big{)}}{\sigma\big{(}1+d(e^{\theta}-1)\big{)}}&0\\ 0&\frac{de^{\frac{-i\pi(\alpha+\beta)}{2}}\big{(}e^{\theta}-1\big{)}\sigma}{1+d(e^{\theta}-1)}&\frac{e^{\frac{\theta}{2}}}{1+d(e^{\theta}-1)}&0\\ 0&0&0&\frac{d(1-e^{\theta})+e^{\theta}}{1+d(e^{\theta}-1)}\end{pmatrix}$ (3.19)
and it reproduces
pure RR / Fendley-Intriligatorinterpolationmixed flux
[…]
$d\to\frac{1}{2},\,\,(\alpha,\beta)=(\frac{1}{2},\frac{1}{2}),\,\,\sigma=-1,\,\,\mbox{for F.-I. or}\qquad\qquad\sigma=-e^{\frac{i\pi(\alpha+\beta)}{2}-\frac{i\pi}{k}},\,\,d\to\frac{1}{2}\big{(}1-i\mbox{cot}\frac{\pi}{k}\big{)}.$
$d\to\frac{1}{2},\,\,(\alpha,\beta)=(0,0),\,\,\sigma=-1,\,\mbox{for \mbox{$AdS_{3}$}}$ (3.20)
We have summarised the situation in fig. 4. Notice that also $S_{a+}$ satisfies braiding unitarity provided $\Phi(\theta)\Phi(-\theta)=1$. It is also physical-unitary if $\Phi(\theta)$ is a pure phase for real $\theta$, and if a certain condition holds on $d$, specifically
$\mbox{Re}(d)=\frac{1}{2},$ (3.21)
@@
(satisfied at least by all the ends of the diagram 3.20) and if $\alpha$ and $\beta$ are real.
Figure 4: $S_{a+}$ interpolates between pure RR massless $AdS_{3}$, Fendley-Intrilgator, and relativistic mixed flux.

### v4

@@ III Relaxing the special coupling
The basic idea is now to let $q$ be generic in this reformulation of the problem, and see where this leads. The hope is that this introduces the mixed flux in the $AdS_{3}$ case - we mean this purely in terms of functional form of the $S$-matrix, since it is clear that the spectra are different [23] and it is not immediate to us to imagine how they might smoothly deform one into another. We first recalculate the $S$-matrix with generic $q$. Applying this to Fendley-Intriligator of course returns back the ordinary Sine-Gordon. For $AdS_{3}$ the result, although potentially interesting per se, is not immediately connected to mixed-flux scattering: this is due to the fact that the mixed-flux relativistic coproduct is twisted in a very particular way [36, 23]. But there is a way to obtain exactly the mixed flux scattering matrix using this philosophy, by adjusting things in such a way that the deformation of the coproduct is precisely the one observed in the mixed flux case. It works like this. We take the following route to the pure RR massless $AdS_{3}$ $S$-matrix:
$S_{k}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&e^{\frac{\theta}{2}}\big(1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big)&\epsilon(\theta)&0\\ 0&\tau_{1}&e^{\frac{\theta}{2}}\big(1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big)&0\\ 0&0&0&\tau_{2}\end{pmatrix}$,
$\tau_{1}=e^{i\frac{\pi}{k}}\big(1-e^{\theta}+e^{i\frac{\pi}{k}+\theta}\epsilon(\theta)\big),\qquad\tau_{2}=e^{\theta}-e^{i\frac{\pi}{k}}\big(1+e^{\theta}\big)\epsilon(\theta)),$
[…]
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}e^{-i\pi(q+i)}\otimes\hat{F}$,
$\hskip 18.49988pt\hskip 18.49988pt|\phi|=|\psi|=0,\qquad|\hat{E}|=|\hat{F}|=0,\qquad q_{0}=-i,$
$\hskip 18.49988pt\hskip 18.49988ptq=\frac{1}{k}-i,\qquad\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi},\qquad\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi},$
@@
$\epsilon(\theta)=\frac{1}{e^{\frac{i\pi}{k}}+\frac{1}{e^{g}-e^{g+\theta}}},\hskip 18.49988ptg\,\,\mbox{constant}.$ (22)
We have solved YBE by noticing that it consistently reduces to a single condition for the function $\epsilon(\theta)$ with various arguments. That is to say that every entry of the Yang-Baxter equation are in fact all one and the same scalar equation for the single function $\epsilon(\theta)$. The form of this equation is $F\Big(\epsilon(\theta_{1}-\theta_{2}),\epsilon(\theta_{1}-\theta_{3}),\epsilon(\theta_{2}-\theta_{3}),...\Big)=0$, with $F$ a certain scalar algebraic expression (the dots denote the explicit $\theta_{i}$ and $k$ dependence coming from the other parts of the $S$-matrix). Having ascertained this fact by brute force, the rest is not difficult: essentially one can for instance solve for $\epsilon(\theta_{1}-\theta_{2})$, and impose the result to be independent of $\theta_{3}$ (by taking the derivative w.r.t. $\theta_{3}$ and setting it to zero). This can be repeated for all the variables. In the resulting equations, one can then set in turn each of the $\theta_{i}$ to zero and permute the arguments. Some simple manipulations ultimately allow one to reduce to an ordinary differential equation for the function $\epsilon(\theta)$ alone: one can then solve this differential equation, and one obtains the solution (22).
@@
$\lim_{\begin{subarray}{c}k\to\infty\\ \epsilon(\theta)\to\tanh\frac{\theta}{2}\end{subarray}}S_{k}\,\,=\,\,\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&\tanh\frac{\theta}{2}&0\\ 0&-\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$, (23)
$\Delta{\hat{E}}\to\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}\otimes\hat{E},\hskip 18.49988pt\Delta{\hat{F}}\to\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}\otimes\hat{F}$.
The choice displayed in the limit (23) consistently corresponds to setting $k\to\infty$, and then setting $e^{g}=-\frac{1}{2}$, in (22).
@@
$S_{k}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&e^{\frac{\theta}{2}}\big(1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big)&\epsilon(\theta)&0\\ 0&\tau_{1}&e^{\frac{\theta}{2}}\big(1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big)&0\\ 0&0&0&\tau_{2}\end{pmatrix}$,
$\tau_{1}=e^{i\frac{\pi}{k}}\big(1-e^{\theta}+e^{i\frac{\pi}{k}+\theta}\epsilon(\theta)\big),\qquad\tau_{2}=e^{\theta}-e^{i\frac{\pi}{k}}\big(1+e^{\theta}\big)\epsilon(\theta)),$
[…]
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}e^{-i\pi(q+i)}\otimes\hat{F}$,
$\hskip 18.49988pt\hskip 18.49988pt|\phi|=|\psi|=0,\qquad|\hat{E}|=|\hat{F}|=0,\qquad q_{0}=-i,\hskip 18.49988pt$
$\hskip 18.49988pt\hskip 18.49988ptq=\frac{1}{k}-i,\qquad\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi},\qquad\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi},$
$\lim\limits_{\epsilon(\theta)\to-i\frac{\sinh\frac{\theta}{2}}{\sin\big(\frac{\pi}{k}-\frac{i\theta}{2}\big.)}}S_{k}\,\,\,=\,\,\,\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt$
$\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{i\sin\frac{\pi}{k}}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}&\frac{\sinh\frac{\theta}{2}}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}&0\\ 0&\frac{\sinh\frac{\theta}{2}}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}&\frac{i\sin\frac{\pi}{k}}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}&0\\ 0&0&0&\frac{\sinh\big(\frac{i\pi}{k}-\frac{\theta}{2}\big)}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}\end{pmatrix}$. (24)
The limit coincides with the left-left relativistic mixed-flux scattering - see for instance [23]. The choice displayed in the limit (24) consistently corresponds to setting $e^{g}=\frac{i}{2}\mbox{csc}\frac{\pi}{k}$, in (22).
@@
$S_{k}=\Phi(\theta)\times$
$\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{\theta}{2}}}{1-e^{i\frac{\pi}{k}+g}\big(e^{\theta}-1\big)}&\frac{1}{e^{\frac{i\pi}{k}}+\frac{1}{e^{g}-e^{g+\theta}}}&0\\ 0&\frac{e^{i\frac{\pi}{k}}\big(1-e^{\theta}\big)\big(1+e^{i\frac{\pi}{k}+g}\big)}{1-e^{i\frac{\pi}{k}+g}\big(e^{\theta}-1\big)}&\frac{e^{\frac{\theta}{2}}}{1-e^{i\frac{\pi}{k}+g}\big(e^{\theta}-1\big)}&0\\ 0&0&0&\frac{e^{\theta}+e^{i\frac{\pi}{k}+g}\big(e^{\theta}-1\big)}{1-e^{i\frac{\pi}{k}+g}\big(e^{\theta}-1\big)}\end{pmatrix}$,
Notice also that $S_{k}$ satisfies braiding unitarity provided $\Phi(\theta)\Phi(-\theta)=1$. It is also physical-unitary if $\Phi(\theta)$ is a pure phase for real $\theta$ and real $k$, and if a certain condition holds on $g$, specifically
$\mbox{Im}(g)=-i\log[-\frac{1}{2}e^{-\mbox{Re}(g)-\frac{i\pi}{k}}\Big(1+\sqrt{1-4e^{2\mbox{Re}(g)}}\Big)\Big],$
In mathematical terms one can introduce a Drinfeld twist to transform the $R$-matrices, for example
$\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}e^{i\pi(q+i)}\otimes\hat{E}=T(\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}\otimes\hat{E})T^{-1},$
$\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}e^{-i\pi(q+i)}\otimes\hat{F}=T(\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}\otimes\hat{F})T^{-1}$
$\hskip 18.49988ptT=e^{i\frac{\pi(q+i)}{2}\mathbbmss{1}\otimes\sigma_{3}},\hskip 18.49988ptq=\frac{1}{k}-i,$ (25)
where $\sigma_{3}=E_{11}-E_{22}$. From the general theory of Drinfeld twists one therefore obtains the relation
$R_{k}=PS_{k},\hskip 18.49988ptP=E_{ij}\otimes E_{ji},\hskip 18.49988ptT_{21}^{-1}R_{k}T=\bar{R}.$ (26)
Similar considerations are at the basis of [57], see also [4] and section 4.3 of [33] for example. Notice that one can use either the $S$-matrix or the associated $R$-matrix - we have actually used the appropriate $R$-matrices throughout the paper whenever we needed to prove that the Yang-Baxter equation
$R_{12}(\theta_{1}-\theta_{2})R_{13}(\theta_{1}-\theta_{3})R_{23}(\theta_{2}-\theta_{3})=$
$\hskip 18.49988pt\hskip 18.49988ptR_{23}(\theta_{2}-\theta_{3})R_{13}(\theta_{1}-\theta_{3})R_{12}(\theta_{1}-\theta_{2})$ (27)
is satisfied in the tensor product space of three particles. The connection with the theory of Drinfeld twists suggests that some of the deformations which we are presenting in this paper are associated with a change in the integrable boundary conditions in the sigma-model picture. One cannot speak of a spin-chain properly in the case of massless scattering [14], but rather one should interpret this as an $S$-matrix encoding a continuum model near criticality [44].
III.1 Fractional route
It is again interesting to obtain an $S$-matrix which depends on a (more general) unspecified fractional fermionic number in such a way to interpolate between Sine-Gordon and $AdS_{3}$ when q=iq=-i, but then to obtain the mixed flux $S$-matrix in another regime.
More general fractional interpolation:
$S_{\alpha+}=\Phi(\theta)\times$
$\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{i\pi(\alpha+\beta)}{2}}\big(-1+e^{\frac{\theta}{2}}\kappa(\theta)\big)}{\sigma}&\kappa(\theta)&0\\ 0&\kappa(\theta)&\sigma e^{\frac{-i\pi(\alpha+\beta)}{2}}\big(-1+e^{-\frac{\theta}{2}}\kappa(\theta)\big)&0\\ 0&0&0&\tau\end{pmatrix}$,
$\tau=-1+2\kappa(\theta)\cosh\frac{\theta}{2},$ (28)
subject to
$\sigma^{2}=1,\hskip 18.49988pt$ (29)
@@ III.1 Fractional route
$\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+e^{2H(\log[q]-\log[-i])-\mbox{Fracto}}\otimes\hat{F}$,
$|\phi|,|\psi|=0,\qquad|\hat{E}|=0,\qquad|\hat{F}|=0,,$
$\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi},\qquad\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi},\,\,\,\,\mbox{Fracto}=\log(-1)\begin{pmatrix}\alpha&0\\ 0&\beta-1\end{pmatrix}$,
@@
$e^{i\pi\alpha}=-q^{2}\,e^{i\pi\beta},\hskip 18.49988ptq=i\sigma e^{\frac{i\pi}{2}(\alpha-\beta)}.$ (30)
@@
$S_{\alpha+}\to\mbox{Fendley-Intriligator if}\,$
$q\to-i,\,\,\alpha,\beta\to\frac{1}{2},\qquad\kappa(\theta)=\mbox{sech}\frac{\theta}{2},\hskip 18.49988pt\sigma=-1,$ (31)
@@
$S_{\alpha+}\to\mbox{massless $AdS_{3}$ if}\,$
$q\to-i,\,\,\alpha,\beta\to 0,\qquad\kappa(\theta)=\mbox{sech}\frac{\theta}{2},\hskip 18.49988pt\sigma=-1.$ (32)
@@
$\sigma=-e^{\frac{i\pi(\alpha+\beta)}{2}-\frac{i\pi}{k}},\hskip 18.49988pt\kappa(\theta)=\frac{e^{\frac{\theta}{2}}\big(e^{\frac{2i\pi}{k}}\ -1\big)}{e^{\frac{2i\pi}{k}+\theta}-1},$ (33)
@@
$S_{\alpha+}\to\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{i\sin\frac{\pi}{k}}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}&\frac{\sinh\frac{\theta}{2}}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}&0\\ 0&\frac{\sinh\frac{\theta}{2}}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}&\frac{i\sin\frac{\pi}{k}}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}&0\\ 0&0&0&\frac{\sinh\big(\frac{i\pi}{k}-\frac{\theta}{2}\big)}{\sinh\big(\frac{i\pi}{k}+\frac{\theta}{2}\big)}\end{pmatrix}$,
which is exactly the mixed-flux result. Given that $\sigma$ is a sign, we must have
$\beta=\frac{1}{k}-\alpha-\frac{2i\log(-\sigma)}{\pi}+4n=$
@@
$\alpha=\beta=(\frac{1}{k}+2n+1,\frac{1}{k}+2n),\hskip 18.49988pt\hskip 18.49988ptn\in\mathbbmss${Z}, (34)
@@
$|\phi|=\frac{1}{k}+1,\hskip 18.49988pt|\psi|=\frac{1}{k},$ (35)
@@
$|\phi|=\frac{1}{k},\hskip 18.49988pt|\psi|=\frac{1}{k}-1.$ (36)
The two choices (35) and (36) are in fact mathematically isomorphic, being simply related to each other by a $\mathbb{Z}_{2}$ symmetry which swaps the two basis elements combined with a $2\pi$ argument-shift, and they are therefore physically equivalent.
@@
$\kappa(\theta)=\frac{e^{\frac{\theta}{2}}}{1+2\,d\,e^{\frac{\theta}{2}}\sinh\frac{\theta}{2}},\hskip 18.49988ptd\,\ \mbox{constant}.$ (37)
@@
$S_{\alpha+}=\Phi(\theta)\times$
$\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{\theta}{2}}}{1+d(e^{\theta}-1)}&\frac{(1-d)e^{\frac{i\pi(\alpha+\beta)}{2}}\big(e^{\theta}-1\big)}{\sigma\big(1+d(e^{\theta}-1)\big)}&0\\ 0&\frac{de^{\frac{-i\pi(\alpha+\beta)}{2}}\big(e^{\theta}-1\big)\sigma}{1+d(e^{\theta}-1)}&\frac{e^{\frac{\theta}{2}}}{1+d(e^{\theta}-1)}&0\\ 0&0&0&\frac{d(1-e^{\theta})+e^{\theta}}{1+d(e^{\theta}-1)}\end{pmatrix}$,
Notice that also $S_{\alpha+}$ satisfies braiding unitarity provided $\Phi(\theta)\Phi(-\theta)=1$. It is also physical-unitary if $\Phi(\theta)$ is a pure phase for real $\theta$, and if a certain condition holds on $d$, specifically
$\mbox{Re}(d)=\frac{1}{2},$ (38)
@@
and if $\alpha$ and $\beta$ are real.